What color bands would be printed on resistors with the resistances below?

(i)    530 kΩ

band 1    
band 2    
band 3    


(ii)    8.4 kΩ

band 1    
band 2    
band 3    
Answer #1

color codes are as follows:

Black - 0

Brown - 1

red - 2

orange - 3

yellow - 4

green -5

blue -6

violet -7

grey -8

white-9

part 1:

530 kilo ohms=53*10^4 ohms

band 1- green (for 5)
band 2- orange (for 3)
band 3-yellow (for exponent 4)


part 2:

8.4 kilo ohms=8.4*10^3=84*10^2

band 1- grey (for 8)
band 2- yellow (for 4)
band 3-red (for exponent 2)